OpenStudy (anonymous):

2x^2 - 13x + 15 = 0?

OpenStudy (saifoo.khan):

do u have like a million questions?

OpenStudy (anonymous):

lol yes:/

OpenStudy (anonymous):

i missed school for a week ha

OpenStudy (saifoo.khan):

try this one!! i can't give u direct answers.. it's a crime on this site. :(

OpenStudy (anonymous):

im doing otheres!

OpenStudy (saifoo.khan):

try it. :) just try.

OpenStudy (anonymous):

im not good at math

OpenStudy (anonymous):

whats 15? its 5x3=15, try (2x-5)(x-5)=0, but here -5x-10x=-15 try (x-5)(2x-3)=0, -10x-3x=-13 ok therefore the answer is... (x-5)(2x-3)=0, x-5=0...x=5 ans..... 2x-3=0,...x=3/2 ans...

OpenStudy (anonymous):

even though you are not good in math, try solving and remember practice makes perfect...lol

OpenStudy (anonymous):

have fun solving...lol